Please Do number 6. the R value for number 5 is 1.23

Please Do number 6.

the R value for number 5 is 1.23
Answer #1

4) If there is a convection heat transfer on either sides of the wall then thre are 5 resistance to heat flow through the window. For unit area (A=1)of window they are,

  • Inside convection resistance R1 = 1/hinside
  • Wall conduction resistance R2= L1/k2
  • Convection resistance in between the walls R3 = 1/hbetween
  • Conduction resistance of wall R4 = L2/k2
  • Outside convection resistance R5 = 1/houtside

So draw 5 resistance in series connection5) R = 1 / U = 1 / 0.81 = 1.23 K/W

when convection on both side is included R value increases

6) Neglecting inside and outside convection to get the published U - factor the convection coefficient must be h = U = 0.81 W/m2k only .